Could you please tell me if I'm on the right track with this equation? This is how I solved it but I'm not sure. Thank you
3x+y-4=0
x=-1/3y+7/3
3(-1/3y+7/3)+y-4=0
-1y+7+1y-4=0
3=0
No solution

Answer by fractalier(6550)

The ones aren't necessary but
3x+y-4=0
x=(-1/3)y+7/3
3((-1/3)y+7/3)+y-4=0
-y+7+y-4=0
3=0
looks correct to me.
That means that these lines, once graphed, are parallel.